OpenStudy (anonymous):

-4+3x+9/7x^3+7-x+7/4x^3

OpenStudy (anonymous):

help plz

OpenStudy (anonymous):

\[ -4+3x+\frac{9x^3}{7}+7-x+\frac{7x^3}{4} \] \[ 3+2x+\frac{36x^3+49x^3}{28} \] \[ 3+2x+\frac{85x^3}{28} \]

OpenStudy (anonymous):

thanks